# JWST’s little red dots: an emerging population of young, low-mass AGN cocooned in dense ionized gas

### v1

@@ Abstract
JWST has uncovered large numbers of compact galaxies at high redshift with broad hydrogen/helium lines. These include the enigmatic population known as “little red dots” (LRDs)[1, 2]. Their nature is debated, but they are thought to be powered by supermassive black holes (SMBHs)[3, 4] or intense star formation[5]. They exhibit unusual properties for SMBHs, such as black holes that are overmassive for their host galaxies[4] and extremely weak X-ray[6, 7, 8, 9] and radio[10, 11, 12] emission. Using the highest-quality JWST spectra, we show here that the lines are broadened by electron scattering with a narrow intrinsic line core. The data require high electron column densities and compact sizes (light days), which, when coupled with their high luminosities can only be explained by SMBH accretion. The narrow intrinsic cores of the lines imply upper limits on the black hole masses of M⊙, two orders of magnitude lower than previous estimates. These are among the lowest mass SMBHs known at high redshift and suggest that this is a population of young, rapidly growing SMBHs. They are enshrouded in a dense cocoon of ionized gas, probably related to their youth, from which they are accreting close to the Eddington limit. Reprocessed nebular emission from the dense cocoon dominates the optical spectrum, explaining most LRD spectral characteristics and helping to suppress radio and X-ray emission.[13, 14]
@@
To test different broad line shapes, we constructed a sample of all broad line galaxies (H km s-1) with high signal-to-noise ratio (SNR) JWST/NIRSpec medium-resolution () spectra [17] in the DAWN JWST Archive (DJA)[18]. Details of the sample and selection criteria are presented in the Methods and Extended Data Table 2. Our search yielded 12 objects at and 18 additional objects for a combined ‘stacked’ spectrum at . High-resolution () data were used where available (Objects A and C) and gave consistent results but better constrained narrow component widths.
@@
As mentioned above, the AGN interpretation of the broad line objects and LRDs discovered by JWST faces several challenges: the large inferred black hole masses, both compared to their host galaxies, and as a function of the age of the universe and their cosmic number density, and the lack of X-ray and radio detections. First, as the electron scattering dominates the line widths in these systems, we need to use the intrinsic line widths prior to scattering, to infer the black hole masses, assuming these intrinsic line widths are entirely due to Doppler orbital motions around the SMBH. These are an order of magnitude lower than inferred with a simple Gaussian fit to the full line profile. The inferred black hole masses, using standard single-epoch black hole mass scaling relations[35], are therefore lower by about a factor of a hundred (Table 1). Our estimates of the maximum SMBH masses are consistent with the typical host galaxy stellar mass found at lower redshifts[35], as shown in Fig. 4, alleviating much of this problem. The much lower black hole masses also alleviate the number density and early formation and growth problems. Second, Compton-thick ionized gas will hinder X-ray[26] and suppress radio[27] emission, by photoelectric and free-free absorption[36] respectively. This is true in X-rays even at low metallicity due to He absorption[37]. The ionized column densities we infer here will attenuate the X-rays by only a factor of a few (see Methods)–less than the dex required to satisfy the observations[6, 7, 8]. Steep hard X-ray spectral slopes and/or relatively low energy power-law cut-offs are also required to suppress the X-rays[6], both of which are found in narrow-line Seyfert 1 objects[38]. Narrow-line Seyfert 1s are believed to be low-mass, high accretion rate SMBHs[39].
@@
The high density gas reprocesses essentially all of the Lyman continuum radiation from the central source, resulting in extremely high Lyman optical depths that cause the self-reversal seen as absorption in the line centres of some of the Balmer lines. The bulk of the ionizing flux from the AGN is therefore emitted in the Balmer and Paschen regions of the spectrum, implying that nebular gas emission must be a major component of the optical and NIR spectra of these sources, giving rise to Balmer, Paschen and He i emission lines and continua and the two-photon continua that would dictate the spectral shape[13, 40]. In addition, while many of these systems appear in blue star-forming galaxies based on the spatially-extended nebulosity surrounding the point source[41], many may be entirely dominated by the nebular emission at optical and even near UV wavelengths. We also note that our inferred electron-scattering optical depths are all (see Extended Data Fig. 8). This suggests three things. First, that there may be a population that is even more heavily obscured, suppressing line emission due to self-absorption, and giving rise to spectra that appear in absorption rather than emission, including a Balmer break instead of a jump. Second, it also suggests that the gas distribution is close to spherical without a large opening, because scattered light will be preferentially emitted along the lowest column density sightlines. The fact that we are not detecting large numbers of low-mass, rapidly accreting AGN without these signatures indicates that LRDs are the primary population and that the gas cocoon in these SMBHs is a near-universal, young phase of their accretion-powered growth. Third, that the line ratios of the Balmer lines should be very different from the standard case B values due to very high Lyman optical depths [42], so that high H/H ratios do not imply dust extinction.
| ID | Field PID-SRCID | Redshift | MBH | M | M⋆ | ||
@@
| A | GN 1181-68797 | 5.0405 | |||||
| B | GN 2674-14 | 5.1826 | |||||
| C | GN 1181-73488 | 4.1327 | |||||
| D | EGS 4233-42046 | 5.2757 | |||||
| E | EGS 4233-49140 | 6.6847 | |||||
| F | EGS 1345-1244 | 4.4771 | — | ||||
| G | EGS 4233-58237 | 3.6505 | |||||
| H | EGS 4106-51623 | 4.9511 | |||||
| I | GN 1181-53501 | 3.4294 | |||||
| J | GN 1181-38147 | 5.8694 | |||||
| K | EGS 4233-50052 | 5.2392 | |||||
| La | EGS 4233-60935 | 5.2877 | |||||
@@
All of this points to a solution to JWST’s riddle of LRDs/compact broad-line objects. They are intrinsically narrow-line AGN, i.e. young[44], low-mass (– M⊙), Eddington-accreting SMBHs[39], buried in a thick cocoon of gas, presumably related to their early stage of formation[45]. Their high accretion rates produce copious UV emission that ionizes their gas cocoon and at the same time efficiently cool and weaken the corona, suppressing their hard X-rays[39]. The ionized gas hinders the escape of radio and X-rays, while reprocessing almost all the Lyman radiation into nebular optical emission, producing the broadened Balmer lines and continuum breaks that characterise the classic ‘v’-shape spectrum[20]. Many galaxies have been found with stellar masses and ages that are so massive and old that they are difficult to explain in standard scenarios[46, 47]. Our results and other recent work[13, 14] indicate that galaxies with red colours and even those with Balmer breaks, may, instead of being massive, quiescent or post-starburst galaxies, be low-mass AGN systems with extreme optical depth ionized gas. The distribution of electron column densities and the lack of a similar or larger population of low-mass, high accretion AGN without the nebular cocoon hints that we may be observing the bulk growth of SMBHs in the phase when they are surrounded by a quasi-spherical, dense gas shell, before angular momentum transport has flattened their accretion into a well-defined disk.

### v2

@@ Abstract
JWST has uncovered large numbers of compact galaxies at high redshift with broad hydrogen/helium lines. These include the enigmatic population known as “little red dots” (LRDs)[1, 2]. Their nature is debated, but they are thought to be powered by supermassive black holes (SMBHs)[3, 4] or intense star formation[5]. They exhibit unusual properties for SMBHs, such as black holes that are overmassive for their host galaxies[4] and extremely weak X-ray[6, 7, 8, 9, 10] and radio[6, 11, 12, 13] emission. Using the highest-quality JWST spectra, we show here that the lines are broadened by electron scattering with a narrow intrinsic line core. The data require high electron column densities and compact sizes (light days), which, when coupled with their high luminosities can only be explained by SMBH accretion. The narrow intrinsic cores of the lines imply upper limits on the black hole masses of M⊙, two orders of magnitude lower than previous estimates. These are among the lowest mass SMBHs known at high redshift and suggest that this is a population of young, rapidly growing SMBHs. They are enshrouded in a dense cocoon of ionized gas, probably related to their youth, from which they are accreting close to the Eddington limit. Reprocessed nebular emission from the dense cocoon dominates the optical spectrum, explaining most LRD spectral characteristics and helping to suppress radio and X-ray emission.[14, 15]
@@
To test different broad line shapes, we constructed a sample of all broad line galaxies (H km s-1) with high signal-to-noise ratio (SNR) JWST/NIRSpec medium-resolution () spectra [18] in the DAWN JWST Archive (DJA)[19, 20]. Details of the sample and selection criteria are presented in the Methods and Extended Data Table 2. Our search yielded 12 objects at and 18 additional objects for a combined ‘stacked’ spectrum at . High-resolution () data were used where available (Objects A and C) and gave consistent results but better constrained narrow component widths.
@@
As mentioned above, the AGN interpretation of the broad line objects and LRDs discovered by JWST faces several challenges: the large inferred black hole masses, both compared to their host galaxies, and as a function of the age of the universe and their cosmic number density, and the lack of X-ray and radio detections. First, as the electron scattering dominates the line widths in these systems, we need to use the intrinsic line widths prior to scattering, to infer the black hole masses, assuming these intrinsic line widths are entirely due to Doppler orbital motions around the SMBH. These are an order of magnitude lower than inferred with a simple Gaussian fit to the full line profile. The inferred black hole masses, using standard single-epoch black hole mass scaling relations[37], are therefore lower by about a factor of a hundred (Table 1). Our estimates of the maximum SMBH masses are consistent with the typical host galaxy stellar mass found at lower redshifts[37], as shown in Fig. 4, alleviating much of this problem. The much lower black hole masses also alleviate the number density and early formation and growth problems. Second, Compton-thick ionized gas will hinder X-ray[28] and suppress radio[29] emission, by photoelectric and free-free absorption[38] respectively. This is true in X-rays even at low metallicity due to He absorption[39]. The ionized column densities we infer here will attenuate the X-rays by only a factor of a few (see Methods)–less than the dex required to satisfy the observations[6, 8, 7, 10]. Steep hard X-ray spectral slopes and/or relatively low energy power-law cut-offs are also required to suppress the X-rays[10], both of which are found in narrow-line Seyfert 1 objects[40]. Narrow-line Seyfert 1s are believed to be low-mass, high accretion rate SMBHs[41].
@@
The high density gas reprocesses essentially all of the Lyman continuum radiation from the central source, resulting in extremely high Lyman optical depths that cause the self-reversal seen as absorption in the line centres of some of the Balmer lines. The bulk of the ionizing flux from the AGN is therefore emitted in the Balmer and Paschen regions of the spectrum, implying that nebular gas emission must be a major component of the optical and NIR spectra of these sources, giving rise to Balmer, Paschen and He i emission lines and continua and the two-photon continua that would dictate the spectral shape[14, 43]. In addition, while many of these systems appear in blue star-forming galaxies based on the spatially-extended nebulosity surrounding the point source[44], many may be entirely dominated by the nebular emission at optical and even near UV wavelengths. We also note that our inferred electron-scattering optical depths are all (see Extended Data Fig. 8). This suggests three things. First, that there may be a population that is even more heavily obscured, suppressing line emission due to self-absorption, and giving rise to spectra that appear in absorption rather than emission, including a Balmer break instead of a jump[45, 46]. Second, it also suggests that the gas distribution is close to spherical without a large opening, because scattered light will be preferentially emitted along the lowest column density sightlines. The fact that we are not detecting large numbers of low-mass, rapidly accreting AGN without these signatures indicates that LRDs are the primary population and that the gas cocoon in these SMBHs is a near-universal, young phase of their accretion-powered growth. Third, that the line ratios of the Balmer lines should be very different from the standard case B values due to very high Lyman optical depths [47], so that high H/H ratios do not imply dust extinction.
| ID | Survey-Field-MSAID | Redshift | MBH | M | M⋆ | ||
@@
| A | JADES-GN-68797 | 5.0405 | |||||
| B | 2674-GN-14 | 5.1826 | |||||
| C | JADES-GN-73488 | 4.1327 | |||||
| D | RUBIES-EGS-42046 | 5.2757 | |||||
| E | RUBIES-EGS-49140 | 6.6847 | |||||
| F | CEERS-EGS-1244 | 4.4771 | — | ||||
| G | RUBIES-EGS-58237 | 3.6505 | |||||
| H | 4106-EGS-51623 | 4.9511 | |||||
| I | JADES-GN-53501 | 3.4294 | |||||
| J | JADES-GN-38147 | 5.8694 | |||||
| K | RUBIES-EGS-50052 | 5.2392 | |||||
| La | RUBIES-EGS-60935 | 5.2877 | |||||
@@
All of this points to a solution to JWST’s riddle of LRDs/compact broad-line objects. They are intrinsically narrow-line AGN, i.e. young[49], low-mass (– M⊙), Eddington-accreting SMBHs[41], buried in a thick cocoon of gas, presumably related to their early stage of formation[50]. Their high accretion rates produce copious UV emission that ionizes their gas cocoon and at the same time efficiently cool and weaken the corona, suppressing their hard X-rays[41]. The ionized gas hinders the escape of radio and X-rays, while reprocessing almost all the Lyman radiation into nebular optical emission, producing the broadened Balmer lines and continuum breaks that characterise the classic ‘v’-shape spectrum[22]. Many galaxies have been found with stellar masses and ages that are so massive and old that they are difficult to explain in standard scenarios[51, 6, 52]. Our results and other recent work[14, 15] indicate that galaxies with red colours and even those with Balmer breaks, may, instead of being massive, quiescent or post-starburst galaxies, be low-mass AGN systems with extreme optical depth ionized gas. The distribution of electron column densities and the lack of a similar or larger population of low-mass, high accretion AGN without the nebular cocoon hints that we may be observing the bulk growth of SMBHs in the phase when they are surrounded by a quasi-spherical, dense gas shell, before angular momentum transport has flattened their accretion into a well-defined disk.